OpenStudy (anonymous):

What is the value of Fraction 1 over 2x3 + 3.4y when x = 3 and y = 4?

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{{2{x^3} + 3.4y}} \to \frac{1}{{2 \times \left( {{2^3}} \right) + \left( {3.4 \times 4} \right)}} = ...?\]

OpenStudy (anonymous):

25.6?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: \[\frac{1}{{2 \times \left( {{2^3}} \right) + \left( {3.4 \times 4} \right)}} = \frac{1}{{16 + 13.6}} = \frac{1}{{29.6}} = ...?\]

OpenStudy (anonymous):

@Michele_Laino im really bad at fractions but i think its 31.6?

OpenStudy (michele_laino):

it is: \[\Large \frac{1}{{2 \times \left( {{2^3}} \right) + \left( {3.4 \times 4} \right)}} = \frac{1}{{16 + 13.6}} = \frac{1}{{29.6}} = 0.0338\] approximated value

OpenStudy (michele_laino):

\[\large \frac{1}{{2 \times \left( {{2^3}} \right) + \left( {3.4 \times 4} \right)}} = \frac{1}{{16 + 13.6}} = \frac{1}{{29.6}} = 0.0338\]

OpenStudy (anonymous):

these are the choices though 15.85 18.1 27.1 31.6

OpenStudy (michele_laino):

Sorry I have made an error since we have to substitute x=3 before Nevertheless, my expression above is wrong.

OpenStudy (anonymous):

oh

OpenStudy (michele_laino):

please write your expression using the editor, or make a drawing of your expression

OpenStudy (anonymous):

when x=3 and y=4

OpenStudy (michele_laino):

ok! then we have: \[\frac{1}{2}{x^3} + 3.4y \to \frac{1}{2} \times {3^3} + 3.4 \times 4 = \frac{{27}}{2} + 13.6 = ...?\]

OpenStudy (anonymous):

27.1

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay thank you (: @Michele_Laino

OpenStudy (michele_laino):

:)